# A Spatial-statistical model to analyse historical rutting data

### 08 Jan 2024 v1

@@ 3.4 Model evaluation
Three model fit measures are available; the deviance Information Criterion (DIC) (Spiegelhalter et al. 2002), the Watanabe-Akaike Information criterion (WAIC) (Watanabe and Opper 2010, Gelman et al. 2014) and the log-likelihood (Hubin and Storvik 2016) can be obtained for each of them in order to select the best one. The DIC and WAIC are based on hierarchical modeling generalizations of the Akaike Information Criterion (Akaike 1974), and are widely used in Bayesian statistics to perform model comparison. The values are determined by the model fit and its complexity, and by a penalty applied for overfitting. Models with a high level of complexity and a high goodness-of-fit, and models with the lowest DIC and WAIC should be chosen.
@@ 4 Results and discussion
We fit the models in Section 3.1.4 to the data in Section 2.3. For interpretability, the explanatory variables are standardized by subtracting the mean and dividing by the standard deviation.
@@
The results from the spatial Model 1 and the non-spatial version Model 4 are used to demonstrate the models’ performance. Based on Model evaluation criteria (Table 5), the spatial and non-spatial models with more explanatory variables (Model 1, 3 or 4) are selected as the best model, with the highest goodness-of-fit (lowest DIC and WAIC). While Models 3 or 5 capture the random (and/or spatial) variability of pavement rutting, they do not necessarily explain more of the uncertainty. Therefore, the spatial and non-spatial Models 1 and 4 are considered the most appropriate model to infer pavement performance from rutting data and are used for further analysis. The results for the remaining models are given in Appendix A.
4.1 Estimated spatial rutting component
Using Model 1 and the estimates obtained from fitting this model, we can identify areas wit h high rutting rates that are not explained by known factors such as pavement type or lane width. This is given by the posterior mean of the spatial rutting component $\omega(s)$. The spatial rutting component $\hat{\omega}(s)$ is plotted in Figure 3.
We find that the rutting inferred from $\hat{\omega}(s)$ is quite variable along the road; graph (a). The mean unexplained rutting contribution from $\hat{\omega}(s)$ is $0.01$ mm/year and the maximum unexplained portion is $3.21$ mm/year. This means that some locations would have $3.21$ mm/year extra rutting compared with the mean. There are also several locations where rutting exceeds $2$ mm/year (the observed mean rutting is $1.67$ mm/year; see Figure 2, graph (a)). This suggests that there are unexplained or unaccounted influential factors of rutting in a given year, one of which could be the weather. Rutting of at least $2$ millimetres at some locations (upper panel), or in a given year (lower panel), is concerning, given the expected lifetime of the pavement and the recommended levels for treatment. If maintenance is to be carried out at $20-25$ millimetres rut depth levels, depending on traffic, and pavement designs are expected to last between $20-30$ years, then some parts of the road would survive at most $12.5$ years ($10$ years in high-traffic areas). The minimum unexplained portion of rutting from $\hat{\omega}(s)$ is $-1.72$ mm/year. This suggests that those locations would have $1.72$ mm/year less rutting compared with mean ($0.01$ mm/year). That is, better real-life rutting performance than expected by the model. The rutting from the spatial component do appear to be relatively systematic, with a difference between the highest and lowest rutting values of $4.93$ millimetres.
@@ 4.1 Estimated spatial rutting component
Including spatial components in the model allow us to explain more of the rutting variability and identify areas with high rutting rates not explained by known variables. This is confirmed in Figure 3 (b), where the rutting explained by the random yearly effect from the spatial model is at least $0.5$ mm/year less compared with the non-spatial model. We further investigate locations with unexpectedly high rutting rates in Section 4.2.
@@
Figure 3: Upper panel: Spatial rutting component (or unexplained portion of rutting $\{\hat{{\omega}}(s)\}$), with approximate 95% credible interval. Lower panel: Rutting from the random yearly effect ($\hat{{\gamma}_{t}}$) from Model 1 ( ) and Model 4 (       ), with approximate 95 % credible interval.
@@ 4.2 Estimated rutting from the explanatory variables, random yearly effect and spatial rutting component
We can also use the spatial model (Model 1) to identify locations, termed ‘hot spots’, that are in the greatest need of maintenance (Figure 4). For illustration, the model is used to infer rutting from the coefficients of the explanatory variables $\hat{\beta}$; the mean of the posterior means of the random yearly effects $\hat{\gamma}$; and the posterior mean of the spatial rutting component $\hat{\omega}(s)$. This is both the explained and unexplained rutting. The random yearly effect $\gamma_{t}$ varies from year to year, so we use an estimate of the mean, i.e., $\hat{\gamma}$. In addition, $\hat{\sigma}_{\xi_{t}}$ varies annually, but it is quite small (see Supplementary Materials LABEL:supplementary:addional_results) so it is not considered in our illustration.
First, we estimate the rutting from the sum of the coefficients of the explanatory variables and the mean of the posterior means of the random yearly effects $\hat{\gamma}$ only. The results are from Model 1, so we denote this by: $\hat{\eta^{*}}=\hat{\beta}_{(1)}z_{(1)(\mathrm{s,t})}+\hat{\gamma}$. Next, we estimate the rutting from the sum of the coefficients of the explanatory variables, the mean of the random yearly effect $\hat{\gamma}$, and posterior mean of the spatial rutting component $\hat{\omega}(s)$; we denote by $\hat{\eta}=\hat{\beta}_{(1)}z_{(1)(\mathrm{s,t})}+\hat{\gamma}+\hat{\omega}(s)$. Since the different asphalt types; stone mastic asphalt (Sma), asphalt concrete (Ac) and asphalt gravel concrete (Agc), are laid at specific locations along E14, these are highlighted in various colours along the plot line $\hat{\eta^{*}}$. These results are shown in Figure 4.
The effect of pavement type on rutting (i.e., the line showing the three asphalt types; Sma, Ac and Agc) is much less variable (ranging from $0.96$ to $3.89$ mm/year) compared with the sum of the explanatory variables, the mean of random yearly effect and the spatial rutting component (ranging from $0.03$ to $4.68$ mm/year); see graph (a). This suggests that the explanatory variables and the random yearly effect alone do not capture or realistically represent the variability in the expected rutting. Accounting for both spatial and random yearly effects allow more of the rutting variability to be explained. These can be seen in Table 3 and Table 4 in Appendix A, where $\hat{\sigma}_{\omega}(s)$ explains $0.47$ and $\hat{\gamma}$ explains $0.72$ compared with the remaining uncertainty ($\hat{\sigma}_{\varepsilon}=$ $1.38$) from the spatial model and the non-spatial version ($\hat{\sigma}_{\varepsilon}=$ 1.54). Additionally, the expected lifetime at some sections of the pavement is at most $6$ years considering only the explanatory variables and the random yearly effect or $5$ years considering the spatial rutting component $\hat{\omega}(s)$. This expected lifetime is depended upon traffic (i.e., $20$ or $25$ millimetres recommended maintenance for high- or low-traffic areas). We also see that the expected rutting $\hat{\eta}$ can be high, where the rutting is explained by the explanatory variables; or it could be moderate, but the spatial rutting component $\hat{\omega}(s)$ is large. Compared with the observed mean rutting across segments, which ranges from $-0.05$ to $8.33$ mm/year (see Figure 2, graph (a)), the maximum expected rutting estimated from the model (i.e., $4.68$ mm/year) is approximately halved the maximum of the observed mean rutting.
In graph (b), the expected rutting from the sum of the explanatory variables, the random yearly effect and the spatial rutting component; lane width; and AADT (per ten thousand) are plotted for a short section of the road. This part of the road, from segments $1500-2700$ (or $30000-54000$ metres towards Sweden), is dominated by the weaker pavement type Agc and, as shown in graph (a), has spikes of accelerated rutting. When plotted against the AADT or lane width, we see that rutting is higher with increasing traffic and narrow road lanes. This large variability may, presumably, be a result of: (i) a change in asphalt type from asphalt concrete (Ac) to asphalt gravel concrete (Agc, see same road stretch in the graph (a)); (ii) the type of traffic in that area, e.g., heavy-traffic (vehicles $>3500$ kilograms); (iii) poor road design; or (iv) lack of timely and efficient pavement rehabilitation.
The observed mean rutting (Figure 2 (a) in Section 2.4) is estimated to be $1.67$ mm/year. From graph (a), we found that the expected rutting, $\hat{\eta}$, for 351 road segments exceed this mean, and 22 of these road segments gave estimated rutting that exceeds $3$ mm/year (graphs (c) and (d), respectively). These 22 locations suggest potential hot spots for accelerated rutting and would require urgent maintenance as the expected lifetime would be at most 8 years, depending on traffic.
(a) Expected rutting from Model 1. Dashed line is mean rutting $1.67$ mm/year.
(b) Expected rutting, AADT and lane width
@@
Figure 4: Graph (a): Expected rutting from: i) the spatial rutting component $\hat{\omega}(s)$, the random yearly effects plus coefficients of all explanatory variables, interacting with AADT for Model 1 $\left\{\hat{\eta}\right\}$ and ii) from the sum of the explanatory variables and the of random yearly effect only $\left\{\hat{\eta^{*}}\right\}$, with each asphalt type highlighted. The effects are inferred for the year 2020. Graph (b): Expected rutting plotted against ADDT and lane width for a small stretch of E14 (closer to Sweden). Graphs (c) and (d) give the locations where rutting exceeds $2$ mm/year or $3$ mm/year, respectively.
4.2.1 Investigating areas with high rutting rates
@@
The large spike in Figure (4) (a), showing high rutting rates, is further investigated. We look at the road segments 2100 - 2500 metres surrounding this spike. This area is the municipality of Meråker, just 20 kilometres west of Storlien. This stretch, as shown in Figure 4 (a), gave high expected rutting, reaching $4.68$ millimetres, with an expected lifetime of just $5$ years (low-traffic area with required maintenance at 25 millimetres rut depth). This accelerated rutting could be because of clogged drainage, flooding, missing side ditches (Figure 5), heavy-traffic or road designs with railguards installed close to the lane edge, allowing that traffic drives more within the same wheel tracks. Meråker has industry and agriculture, therefore, roadways are more subjected to heavy vehicles, including farm vehicles such as tractors (industry, wheel drive, row crop).
(a) Short section between EV14 S6D1 m6148-m8212 F1 with high rutting.
(b) Location EV14 S6D1 m8017 F1 in year 2020
(c) Location EV14 S6D1 m7971 F1 in year 2021
(d) Location EV14 S6D1 m7961 F1 in year 2022
@@
Figure 5: Graph (a): Location with high rutting rates (road segments $2100-2500$ towards Sweden). Graphs (b), (c) and (d) show the location, in years 2020-2022 (see Statens vegvesen or go to the next url: https://vegkart.atlas.vegvesen.no/#kartlag:nib/@337286,7036109,14/vegsystemreferanse:337312.447:7036064.784
To further investigate this road section with high rutting values, we consider the effect of the explanatory variables, random yearly effect, and the spatial rutting component. We explore the spatial Model 1 using various values of the explanatory variables to approximate the rutting. First, we use the true values in the data, i.e., the asphalt type, the lane width and rut depth measured at that location, denoted by $\hat{\eta}_{\mathrm{True}}$. Second, the asphalt type is changed from its true state of asphalt gravel concrete (Agc) to asphalt concrete (Ac) or stone mastic asphalt (Sma); represented by $\hat{\eta}_{\mathrm{Ac}}$ or $\hat{\eta}_{\mathrm{Sma}}$, respectively. Third, the ‘best’ (or ‘worse’) values are used, i.e., if pavement type with the strongest binder, Sma, or a rut depth value from the previous year of $0.01$ millimetres ($\sim 0$), or lane width of $6$ metres were used. This is denoted by $\hat{\eta}_{\mathrm{Best}}$. These values are chosen following the reasoning: ($\sim$) zero rut depth means no rutting; wider roads mean a slower rate of rutting. The values considered to be ‘worse’, resulting in accelerated rutting, include a rut depth from the previous year of 25 millimetres, lane width of 2.75 metres and asphalt gravel concrete (Agc) pavement type. This is denoted by $\hat{\eta}_{\mathrm{Worse}}$. These values are found in the E14 data set. The cases are plotted in Figure 6 (a), which gives the expected rutting from the sum of the explanatory variables, th yearly random effect $\hat{\gamma}$ and the spatial rutting component $\hat{\omega}(s)$.
Using the worse values in the analysis, the expected rutting is at least 2.3 ($2.25-7.94$) mm/year; Figure 6, graph (a) ( ). This means that with low AADT and the recommended maintenance rut depth level of 25 millimetres, pavement deterioration would accelerate and treatment or new pavement would be required within $3$ to $11$ years after initial construction. The best covariate values ( ) gave at most $4.26$ mm/year of rutting, with the least rutting value ($0.21$ mm/year). And, the earliest maintenance activity would be required from around 6 years after construction or maintenance. Further, the best values ( ) gave similar estimates of annual rutting to that of using the stronger asphalt binder type Sma (stone mastic asphalt,  ), except at road segments $2279-2388$, where the estimated rutting from only changing the asphalt type from asphalt gravel concrete to stone mastic is lower than that of using the best values. This suggests that asphalt type is the driving factor of expected rutting. Pavement types asphalt gravel concrete (Agc,  ) and asphalt concrete (Ac,  ) gave similar estimates of expected rutting. This result is expected because they are made with similar binder concentration strengths.

### 14 Feb 2024 v2

@@ 3.5 Model evaluation
Three model fit measures are available; the deviance Information Criterion (DIC) (Spiegelhalter et al. 2002), the Watanabe-Akaike Information criterion (WAIC) (Watanabe and Opper 2010, Gelman et al. 2014) and the log-likelihood (Hubin and Storvik 2016) can be obtained for each of them in order to select the best one. The DIC and WAIC are based on hierarchical modeling generalizations of the Akaike Information Criterion (Akaike 1974), and are widely used in Bayesian statistics to perform model comparison. The values are determined by the model fit and its complexity, and by a penalty applied for overfitting. Models should balance complexity and goodness-of-fit, and models with the lowest DIC and WAIC should be chosen.
@@ 4 Results and discussion
We fit the models in Section 3.2.3 to the data in Section 2.3. For interpretability, the explanatory variables are standardized by subtracting the mean and dividing by the standard deviation. The results from the spatial Model 1 and the non-spatial version Model 4 are used to demonstrate the models’ performance. Based on Model evaluation criteria (Table 5), and for easy comparison of the models, the spatial and non-spatial models with more explanatory variables are selected. The results for the remaining models are given in Appendix A.
4.1 Estimated spatial rutting component and yearly effect
We analyze estimates from Model 1, identifying areas with high rutting rates not explained by AADT, pavement type, and lane width. The spatial rutting component ()$\hat{\omega}(s)$) plotted in Figure 3(a), varies significantly along the road. Some locations exhibit up to $3.15$ mm/year extra rutting beyond what is accounted for by known factors, while others show rutting exceeding $2$ mm/year. Conversely, the minimum unexplained rutting is $-2.45$ mm/year, indicating better performance than expected.
Figure 3(b) presents estimates of the random yearly $\hat{\gamma}_{t}$. Rutting varies notably between years for spatial and non-spatial models, with minimum estimates in 2016 and maximum in 2018. A systematic shift between spatial and non-spatial models is evident, particularly in pavement type coefficients (see Table 3 in Section 4.2).
@@ 4.1 Estimated spatial rutting component and yearly effect
Including spatial components in the model enhances rutting variability explanation and identifies areas with unexplained high rutting rates. This is confirmed in Figure 3 (b), where the rutting explained by the random yearly effect from the spatial model is at least $0.5$ mm/year less compared with the non-spatial model. Further discussion on estimated regression coefficients is in Section 4.2.
@@
Figure 3: Upper panel: Spatial rutting component (unexplained rutting, $\hat{{\omega}}(s)$), with 95% credible interval. Lower panel: Rutting from the random yearly effect ($\hat{{\gamma}_{t}}$) from Model 1 ( ) and Model 4 (       ), with 95 % credible interval.
4.2 Estimated model parameters
The estimated parameters and 95% credible intervals (CI) for spatial (Model 1) and non-spatial (Model 4) are shown in Table 3. Spatial models exhibit notably higher estimates for pavement type coefficients compared with non-spatial models (see also Table A.2 in Appendix A). Moreover, models lacking the spatial rutting component display larger variance in the ‘noise’ term ($\hat{\sigma}{\varepsilon}$), suggesting that the spatial rutting component captures unobservable spatially varying factors such as weather and traffic.
Additionally, the spatial model better explains expected rutting, as indicated by smaller ‘noise’ terms ($\hat{\sigma}{\varepsilon}$) and uncertainty in random year effects ($\hat{\sigma}_{\gamma}$) compared with non-spatial models (refer to Table 3 and Table 4 in Appendix A). Further results are provided in Supplementary Materials LABEL:supplementary:addional_results.
Table 3: Estimated model parameters with $95\%$ credible intervals
Spatial Model Non-spatial Model
Model 1 Model 4
Mean/Median 95% CI Mean/Median 95% CI
$\hat{\beta}_{\mathrm{Ac}}$ 7.09 [6.56, 7.61] 3.27 [2.84, 3.71]
$\hat{\beta}_{\mathrm{Agc}}$ 7.31 [6.08, 7.73] 3.24 [2.78, 3.70]
$\hat{\beta}_{\mathrm{Sma}}$ 3.53 [3.26, 3.80] 2.03 [1.84, 2.22]
$\hat{\beta}_{\mathrm{d-1}}$ 0.76 [0.68, 0.84] 1.07 [0.98, 1.17]
$\hat{\beta}_{\mathrm{w}}$ 0.07 [0.01, 0.13] 0.05 [-0.02, 0.11]
$\hat{\sigma}_{\omega(s)}$ 0.48 [0.45, 0.52] - -
$\hat{\sigma}_{\gamma}$ 0.92 [0.63, 1.42] 1.37 [0.94, 2.10]
$\hat{\sigma}_{\varepsilon}$ 1.38 [1.38, 1.39] 1.54 [1.53, 1.56]
4.2.1 Effects of the explanatory variables
Pavement types, particularly those with weaker binder concentrations like asphalt gravel concrete ($\hat{\beta}_{\mathrm{Agc}}$) and asphalt concrete ($\hat{\beta}_{\mathrm{Ac}}$), experience more rutting (Table 3). Qian et al. (2019) and Mascarenhas et al. (2020) shed light on how temperature and heavy traffic ($>$ 3.5 tonnes) impact rutting resistance on pavement types such as Agc and Ac. Notably, asphalt gravel concrete is predominantly situated approximately 30 kilometers from the Swedish border, where a sub-arctic climate, underdeveloped networks, and heavy traffic converge to adversely affect pavement performance Peel et al. (2007), Ketzler et al. (2021), Jenelius (2010), Hovi et al. (2018), Pinchasik et al. (2020). There is also greater uncertainty in asphalt concrete ($\hat{\beta}_{\mathrm{Ac}}$) and asphalt gravel concrete ($\hat{\beta}_{\mathrm{Agc}}$) compared with stone mastic asphalt ($\hat{\beta}_{\mathrm{Sma}}$), although considering spatial dependencies helps to explain more of the variability. Non-spatial models appear to explain variability between years rather than by pavement type, suggesting a nuanced relationship between these factors in understanding rutting susceptibility.
Rut depth from the previous year ($\hat{\beta}_{\mathrm{d-1}}$) becomes more crucial without accounting for spatial variability (Model 4, Table 3). The estimated effect of lane width ($\hat{\beta}_{\mathrm{w}}$) is minimal and positive, suggesting negligible impact on expected rutting.
These findings indicate that changes in coefficients are significant when spatial dependencies of unmodeled factors (e.g., weather) are considered. Furthermore, the spatial model captures more uncertainty (smaller $\hat{\sigma}_{\varepsilon}$).
@@ 4.3 Estimated rutting from the explanatory variables, random yearly effect and spatial rutting component
We employ Model 1 to identify maintenance ‘hot spots’ where intervention is most critical. This model combines rutting inference from explanatory variable coefficients ${\hat{\beta}_{(\cdot)}}$, the mean of the yearly random effects ($\hat{\gamma}$), and the spatial rutting component ${\hat{\omega}(s)}$. Since the yearly random effects ($\gamma_{t}$) vary annually, we use an estimate of their mean, denoted as $\hat{\gamma}$. The yearly variation in ($\hat{\sigma}_{\xi_{t}}$) is small annually (Supplementary Materials LABEL:supplementary:addional_results), so it is not considered in our analysis. We account for both explained and unexplained rutting.
Initially, rutting estimation relies on explanatory variables and the yearly random effect ($\hat{\eta}^{*}=\hat{\beta}_{{(1)}}z_{(1)(\mathrm{s,t})}+\hat{\gamma}$), with each asphalt type highlighted as Sma ( ), Ac ( ), and Agc ( ) (see Figure 4, graph (a)). Incorporating spatial rutting components yields a more comprehensive estimation ${\hat{\eta}=\hat{\beta}_{(1)}z_{(1)(\mathrm{s,t})}+\hat{\gamma}+\hat{\omega}(s)}$. A comparison of these estimates highlights the influence of spatial components.
The results indicate less variability in rutting estimates without spatial components (ranging from $0.97$ to $3.64$ mm/year) compared to estimates including spatial factors (ranging from $-0.44$ to $4.67$ mm/year). This suggests that explanatory variables and the random yearly effect alone do not fully capture or represent the true rutting variability. Including spatial and yearly effects provides a more comprehensive understanding of rutting variability. Tables 3 and 4 in Appendix A demonstrate this, with the remaining uncertainty ($\hat{\sigma}_{\varepsilon}$) lower in the spatial model ($1.38$ mm/year) compared to the non-spatial model ($1.54$ mm/year).
The findings indicate that high rutting estimates could arise from explanatory variables, while moderate estimates, coupled with large spatial rutting components, suggest the influence of other factors.
(a) $\hat{\eta}^{*}$, $\hat{\eta}$ and mean rutting $1.62$ mm/year (       ).
(b) $\hat{\eta}$, AADT and lane width
@@
Figure 4: Graph (a): Expected rutting from: i) $\hat{\eta}$ and ii) from $*\hat{\eta}{{}^{*}}$, with each asphalt type highlighted. The effects are inferred for the year 2020. Graph (b): Expected rutting plotted against AADT and lane width for a short stretch. Graphs (c) and (d) give the locations where rutting exceeds $2$ mm/year and $3$ mm/year, respectively.
4.3.1 Comparing expected rutting with observed data
The observed mean rutting (Figure 2 (a) in Section 2.4) stands at $1.62$ mm/year. In Figure 4, graph (a), 361 road segments have an expected rutting exceeding this mean, with 49 segments showing estimated rutting exceeding $3$ mm/year (graphs (c) and (d), respectively). These 49 locations signify potential hot spots for accelerated rutting, necessitating urgent maintenance, as their expected lifetime could be as short as 8 years, depending on traffic. Additionally, the observed mean rutting across segments ranges from -0.14 to 8.33 mm/year (Figure 2, graph (a)), while the maximum expected rutting from the model is 4.67 mm/year (Figure 4, graph (a)), approximately half the maximum observed mean rutting.
In Figure 4, graph (b), we plot the expected rutting $\hat{\eta}$, lane width, and AADT (per ten thousand) for a short road section spanning segments $1500-2700$ (or $30-54$ km from Stjørdal), characterized by weaker pavement type Agc. This section exhibits spikes of accelerated rutting, particularly evident with increasing traffic. Possible factors contributing to this variability include changes in asphalt type, traffic composition, suboptimal road design, or inadequate pavement maintenance.
4.4 Investigating an area with high rutting rates
@@
The significant spike in Figure (4) (a), indicating high rutting rates, prompts further investigation. We examine road segments spanning 2100 to 2700 meters surrounding this spike, located in the municipality of Meråker, just 20 kilometers west of the Swedish border at Storlien. This stretch, depicted in Figure 4 (a), exhibits notably high rutting, peaking at $4.67$ mm, with an estimated lifespan of only $5$ years (in a low-traffic area requiring maintenance at 25 mm rut depth). Possible contributors to this accelerated rutting include clogged drainage, flooding, absence of side ditches (see Figure 5), heavy traffic, or road designs featuring railguards installed close to the lane edge, which encourages traffic to remain within the same wheel tracks (refer to Figure A.4 in Appendix A.4). Meråker’s industrial and agricultural activities result in increased traffic from heavy vehicles, including farm equipment such as tractors (industrial, wheel drive, row crop).
(a) Short section between EV14 S6D1 m3841-m4285 F1 with high rutting.
(b) Location EV14 S6D1 m4068 F1 in year 2020
(c) Location EV14 S6D1 m4059 F1 in year 2021
(d) Location EV14 S6D1 m4065 F1 in year 2022
@@
Figure 5: Graph (a): Location with high rutting rates (road segments $2100-2700$ towards Sweden). Graphs (b), (c) and (d) show the location, in years 2020-2022 (see Statens vegvesen or url: https://vegkart.atlas.vegvesen.no/#kartlag:nib/@335074,7038866,15/vegsystemreferanse:335035.067:7038844.352
4.4.1 Exploring the spatial model with various values of the explanatory variables
To further investigate high rutting values in a road section, we analyze the impact of explanatory variables, yearly variation, and spatial rutting component ($\hat{\eta}$). We explore spatial Model 1 using different values for variables to estimate rutting. First, we use actual measurements like asphalt type, lane width, and rut depth ($\hat{\eta}_{\mathrm{True}}$). Second, we consider alternative asphalt types - asphalt concrete (Ac) or stone mastic asphalt (Sma) ($\hat{\eta}_{\mathrm{Ac}}$ or $\hat{\eta}_{\mathrm{Sma}}$). Third, we examine extreme values such as the strongest pavement (Sma), minimal rut depth (0.01 mm), or wider lanes (6 meters) ($\hat{\eta}_{\mathrm{Best}}$). Conversely, accelerated rutting scenarios involve deeper rut depth (25 mm), narrower lanes (2.75 meters), and asphalt gravel concrete (Agc) ($\hat{\eta}_{\mathrm{Worse}}$) from the E14 dataset. These scenarios are depicted in Figure 6 (a).
With worse values ( ), rutting is projected to be at least 1.5 mm/year ($1.76-8.20$), indicating accelerated deterioration and the need for maintenance within 3 to 14 years. Locations with unexpectedly high rutting, exceeding 4 mm/year, are observed, possibly due to increased AADT compared to other areas (Figure 4, graph b).
Using the best covariate values ( ), rutting is projected to be at most 4.30 mm/year, with maintenance required approximately 6 years after construction or maintenance. Additionally, using the stronger asphalt binder Sma yields similar rutting estimates but generally lower than those using the best values, indicating asphalt type as a primary factor. Asphalt gravel concrete (Agc,  ) and asphalt concrete (Ac,  ) show similar rutting estimates due to similar binder concentrations.